# Universal Maximum Likelihood (List) Decoding via Fast Vector-Matrix Multiplication

### v1

@@ Abstract
Maximum-likelihood (ML) decoding for arbitrary block codes remains fundamentally hard, with worst-case time complexity—measured by the total number of multiplications—being no better than straightforward exhaustive search, which requires $q^{k}n$ operations for an $n,k]_{q}$ code. This paper introduces a simple, code-agnostic framework that reduces the worst-case complexity by a factor of $n$, down to $q^{k}$ operations, a highly desirable reduction in practice. The result holds for both linear and nonlinear block codes over general memoryless channels and under both hard-decision and soft-decision decoding. It naturally extends to intersymbol-interference (ISI) channels and ML list decoding with only a negligible increase in complexity. Our core insight is that, upon receipt of each sequence at the receiver, the conditional probability of that sequence for each codeword in the codebook (i.e., the likelihood) can be expressed as the inner product of two carefully constructed vectors—the first depending on the received sequence, and the second on that codeword itself. As a result, evaluating the likelihoods for all codewords in the codebook reduces to a single vector-matrix multiplication, and ML decoding (MLD) becomes the simple task of picking the maximum entry in the resulting vector. The only non-trivial cost lies in the vector-matrix product. However, our matrix construction allows the use of the Mailman algorithm to reduce this cost. This time reduction is achieved at the cost of high space complexity, requiring $\mathcal{O}(q^{k+1}n)$ space to store the pre-computed codebook matrix.
@@ I-A Brief Historical Prospective
Another approach reformulates the ML decoding problem over a discrete memoryless channel symmetric channel (NP-hard in general) as an integer linear program (ILP), whose optimal solution corresponds to the ML codeword; see [15] and related subsequent works. Relaxing the integrality constraints yields a linear program (LP) that can be solved in polynomial time, for instance by the Ellipsoid algorithm [16]. Under certain conditions, the LP relaxation is tight, meaning its solution coincides with that of the original ILP and hence produces the ML codeword in polynomial time. When these conditions fail, however, the LP solution is typically fractional and does not correspond to any valid codeword. In short, this method does not guarantee recovery of the ML codeword in general.
Practical coding standards use particular decoders co-designed with code-specific decoding methods to achieve an ML decoding performance in polynomial time for a restricted class of codes or channels. Prominent examples include Bose–Chaudhuri–Hocquenghem (BCH) codes decoded by the Berlekamp–Massey algorithm [17], decoding of expander codes over the BSC [18], and decoding of first-order Reed-Muller codes [19]. Recently, code-agnostic approaches to ML decoding have attracted significant attention, as they enable universal applicability across arbitrary codes and support practical implementations in reliable communication settings, where the SNR is typically high and moderate-to-high-rate codes are employed. This line of research was initiated by the remarkable work of Duffy et al. [20], who introduced the method of ML decoding via noise guessing, termed GRAND. This framework addresses discrete memoryless additive-noise channels, where the received sequence equals the transmitted codeword plus noise (modulo the alphabet). Because the noise uniquely determines the transmitted codeword given the received sequence, decoding reduces to testing (or, guessing) candidate noise vectors ranked by likelihood. Each is subtracted from the received sequence and checked for codebook membership; the first valid residual yields the ML codeword. The computational cost is governed by the number of guesses required before termination. Although an exact complexity is unknown, it was shown that in the asymptotic regime $n\to\,\infty$, the expected number of guesses scales as $2^{\,n\min\{H_{1/2},\,1-\frac{k}{n}\}}$, where $H_{\alpha}$ denotes the Rényi entropy of order $\alpha$ of the noise (with $H_{1}$ being the Shannon entropy). As a Viterbi decoding approach for linear block codes, this method offers substantially lower complexity than ESD in the high-SNR regime (i.e., when the noise entropy is small) and is particularly effective for high-rate, short block codes. Its advantage, however, vanishes as the SNR decreases, the code rate is reduced, and especially when the channel is non-additive.
@@
As a related note, extensions of ML decoding to list decoding, soft-decision decoding, and erasure decoding have been widely studied. For example, Elias’s early work on list decoding (also called as list-$\ell$ decoding, where $\ell$ denotes the list size) [24] showed that allowing the decoder to output multiple candidates (i.e., allow some level of ambiguity) can reduce error probability, while erasure decoding provides a simple way to incorporate partial reliability information. Soft-decision decoding, particularly with ML-based algorithms, has since become the standard in practice [25]. However, each of these approaches is tailored to a specific setting, and no existing method provides a unified, code-agnostic framework that applies simultaneously to ML hard-decision, soft-decision, erasure, and list decoding.
@@ I-B The Proposed Algorithm and its Relation to the Prior Work
This paper introduces a code-agnostic framework that, for the first time, breaks the long-standing $\mathcal{O}(2^{k}n)$ worst-case complexity barrier of ESD. The key idea is to express the likelihood of each codeword as the inner product of two vectors: one derived from the received sequence, and the other representing the codeword itself. This allows us to represent the entire codebook as a single, large, binary matrix $\mathbf{M}$. Consequently, ML decoding is reduced to a single vector–matrix multiplication, followed by selecting the maximum entry of the resulting vector.
The efficiency of the proposed approach stems from exploiting the discrete structure of $\mathbf{M}$. This performance gain parallels the classical distinction between sorting arbitrary real numbers and sorting integers. In general, sorting $n$ arbitrary items requires $\Omega(n\log n)$ comparisons (i.e., grows on the order of $n\log n$), whereas algorithms such as Counting Sort—particularly efficient when the range of possible input values is small relative to the number of elements—leverage the finite alphabet of integers to achieve linear $\mathcal{O}(n+k)$ complexity [26], where $k$ denotes the range size of the integer-valued inputs. Likewise, rather than treating the entries of $\mathbf{M}$ as arbitrary reals, we exploit its binary structure via the Mailman algorithm [27], which computes the product of an $m\times n$ binary matrix $\mathbf{A}$ with a real vector $\boldsymbol{x}\in\mathbb{R}^{n}$ in $\mathcal{O}(mn/\log\max\{m,n\})$ time—improving upon the straightforward $\mathcal{O}(mn)$ complexity for general matrices. This contrasts with the classical result of Winograd [28], which establishes that general vector-matrix multiplication requires $\Omega(mn)$ operations. Employing this algorithm reduces the total number of operations to $\mathcal{O}(2^{k})$—an $n$-fold improvement over ESD. For the special case of linear codes over the BSC, the framework further adapts to syndrome decoding with a complexity of $\mathcal{O}(2^{\min\{k,\,n-k\}})$, again improving upon the standard $\mathcal{O}(n\cdot 2^{\min\{k,\,n-k\}})$ complexity of trellis-based ML decoders. Since the time complexity is always proportional to the number of codewords (or syndromes, whichever is smaller), no further substantial reduction is possible: identifying the most likely codeword necessarily requires considering each candidate at least once. The main drawback is memory usage: the method requires storing a matrix encoding all codewords in the codebook, which incurs a space complexity of $\mathcal{O}(2^{k+1}n)$. Nevertheless, the framework is universal: it applies to both linear and nonlinear block codes, and supports hard- and soft-decision decoding over general memoryless channels as well as intersymbol-interference (ISI) channels—the most widely studied class of channels with memory [29].
@@
Unlike all previously known ML decoding methods, the proposed decoder universally achieves a complexity of $\mathcal{O}(2^{k})$, which is independent of the block length $n$ when the code rate $k/n$ is assumed to be fixed. This complexity is also independent of the SNR level, distinguishing it from approaches like GRAND [20]. Moreover, the algorithm does not depend on the specific received sequence, unlike the algebraic ML decoder of [31]. The complexity of our proposed method is smaller by a logarithmic factor compared to the best known trellis-based ML decoding via the Viterbi algorithm. This makes the method particularly attractive for low-SNR channels—a relatively uncharted regime—where reliable transmission demands low-rate codes with large minimum distance. In this setting, approximate ML decoding algorithms fail to deliver satisfactory performance, and exact ML decoders cannot achieve lower complexity than ESD. Furthermore, when extended to list decoding, the algorithm’s complexity increases only logarithmically with the list size, in contrast to other approaches such as SOGRAND [32] and List Viterbi [33], whose complexity grows linearly with the list size.
@@
The proposed method is especially relevant for emerging applications such as the ultra-reliable low-latency communication (URLLC), a foundational requirement of 5G and future 6G networks. These systems face a fundamental tradeoff: to increase reliability, one often lengthens the block length $n$, which, for conventional universal decoders, directly increases decoding latency, often scaling as $\mathcal{O}\bigl(2^{k}n\bigr)$. In contrast, our method’s $\mathcal{O}(2^{k})$ complexity decouples latency from block length. There are two key advantages: 1) for a given message/information size $k$, designers can increase $n$ to boost reliability without increasing decoding delay; and 2) the $n$-fold reduction in complexity becomes increasingly significant for large $n$. These benefits position our method as a compelling enabler in domains demanding both high reliability and minimal latency, such as the Tactile Internet, V2X (vehicle-to-everything) communication, and industrial automation [36, 37, 38].
@@
Same time complexity as ESD.
@@
Our method. In general $\mathcal{O}(2^{k})$ $\mathcal{O}(2^{k+1}n)$ Code-agnostic; Works for all codes for both hard- and
        For linear codes $\mathcal{O}(2^{\min\{k,\,n-k\}})$ $\mathcal{O}(2^{\min\{k,\,n-k\}}\cdot 2n)$ soft-decision decoding. High space complexity.
@@ I-C Organization
The core contributions of this paper are presented in Section II. Part II-A introduces the universal decoding method for discrete memoryless channels and analyzes its complexity. Part II-B extends the framework to soft-decision decoding and list-$\ell$ ML decoding, and derives the associated complexities. Part II-C applies the method to erasure decoding. Part II-D focuses on linear codes over symmetric channels, showing how syndrome decoding can be used to further reduce complexity. Finally, Part II-E treats the generalization to $L$-tap ISI channels. We conclude the paper in Section III, where we also discuss the inherent difficulty of the underlying vector-matrix multiplication routine and, drawing on prior work—though without formal proof—argue that further substantial improvements in the worst-case complexity are unlikely. Finally, a brief overview of the Mailman algorithm is given in Appendix The Mailman Algorithm for Fast vector-matrix Multiplication.
@@ II-A Decoding Algorithm Formulation
which can be accomplished using a trivial linear-time scan: by iterating through the vector once and updating a register that stores the current maximum value. Since the length of the resulting vector $\mathbf{m}$ is $S$, this step has a computational complexity of $\mathcal{O}(S)$. Before analyzing the overall decoding complexity, we summarize the procedure in Box II-A and exemplify it in Example 1.
@@
Example 1.
Consider a BSC channel with bit flip probability $p_{e}$ in which the non-linear binary code $\mathscr{C}=\{001,010,100,111\}$ is used. Let $a:=\log P(0\mid 0)=\log(1-p_{e})$ and $b:=\log P(0\mid 1)=\log p_{e}$. Upon receiving a received sequence, say $\boldsymbol{y}=000$, we construct the conditional probability vector as
$V(\boldsymbol{y})=[a,\,b,\,a,\,b,\,a,\,b]\in\mathbb{R}^{6}.$
The codebook matrix $\mathbf{M}$ and the resulting score vector $\mathbf{m}=V(\boldsymbol{y})\cdot\mathbf{M}$ are
$\mathbf{M}=\begin{pmatrix}|&|&|&|\\ I(\boldsymbol{c}^{1})^{\top}&I(\boldsymbol{c}^{2})^{\top}&I(\boldsymbol{c}^{3})^{\top}&I(\boldsymbol{c}^{4})^{\top}\\ |&|&|&|\end{pmatrix}=\begin{pmatrix}1&1&0&0\\ 0&0&1&1\\ \hline\cr 1&0&1&0\\ 0&1&0&1\\ \hline\cr 0&1&1&0\\ 1&0&0&1\end{pmatrix}$,
and
$\mathbf{m}=\begin{pmatrix}2a+b,\,2a+b,\,2a+b,\,3b\end{pmatrix}$.
The ML decision corresponds to the maximum entry in $\mathbf{m}$. When $p_{e}<0.5$, the first three entries, corresponding to the codewords all at Hamming distance 1 from $\boldsymbol{y}$, are maximal. The algorithm requires storing at the receiver a binary matrix $\mathbf{M}$ of size $6\times 4$.
@@
Space complexity. Our algorithm requires constructing and storing a matrix $\mathbf{M}$ that encodes all codewords in the code $\mathscr{C}$, i.e., the entire universe. This matrix has dimensions $nq\times S$. For a $q$-ary linear code of dimension $k$ with $S=q^{k}$ codewords, this results in a space complexity of $\mathcal{O}(q^{k+1}n)$.
@@
Step 2 involves first constructing the matrix $\mathbf{M}$ from all codewords in the code $\mathscr{C}$, followed by its factorization
$\mathbf{M}=UP$
to enable the use of the Mailman algorithm (see Appendix The Mailman Algorithm for Fast vector-matrix Multiplication). Both operations can be performed offline at the receiver and thus are treated as one-time pre-processing steps. Their complexities are amortized and excluded from the complexity analysis. This treatment aligns with standard practices in prior works, such as [34], and is analogous to the pre-processing steps required to construct the trellis in the Viterbi algorithm [9].
Remark 2.
For a $n,k]_{q}$ code, the proposed decoding method has complexity $\mathcal{O}(q^{k})$, which depends only on the alphabet size $q$ and code dimension $k$, but not on the block length $n$. This complexity is smaller by a factor of $n$ (i.e., a logarithmic factor gained through the denominator in Eq. (7)) compared to ESD, as well as other universal decoding methods; see Section I for details, and offers a significant gain for codes with long block length.

### v2

@@ Abstract
Maximum-likelihood (ML) decoding for arbitrary block codes remains fundamentally hard. Its computational complexity—measured by the total number of multiplications—is no better than that of exhaustive search, which requires $q^{k}n$ operations for an $n,k]_{q}$ code. This paper introduces a simple, code-agnostic framework that, for fixed-rate codes (i.e., when $k/n$ remains constant), reduces the complexity of ML decoding by a factor of $n$, achieving $q^{k}$ operations—a highly practical and desirable improvement. The result holds for both linear and nonlinear block codes over general memoryless channels and under both hard-decision and soft-decision decoding. It naturally extends to intersymbol-interference (ISI) channels and list decoding with only a negligible increase in complexity. Our core insight is that, upon receipt of each sequence at the receiver, the conditional probability of that sequence for each codeword in the codebook (i.e., the likelihood) can be expressed as the inner product of two carefully constructed vectors—the first depending on the received sequence, and the second on that codeword itself. As a result, evaluating the likelihoods for all codewords in the codebook reduces to a single vector-matrix multiplication, and ML decoding (MLD) becomes the simple task of picking the maximum entry in the resulting vector. The only non-trivial cost lies in the vector-matrix product. However, our construction allows the use of the Mailman algorithm to reduce this cost. This time reduction is achieved at the cost of higher space complexity, requiring $\mathcal{O}(q^{k})$ space to store the pre-computed matrix containing information of the entire codebook.
@@ I-A Brief Historical Prospective
Another approach reformulates the ML decoding problem over a discrete memoryless channel symmetric channel (-hard in general) as an integer linear program (ILP), whose optimal solution corresponds to the ML codeword; see [17] and related subsequent works. Relaxing the integrality constraints yields a linear program (LP) that can be solved in polynomial time, for instance by the Ellipsoid algorithm [6]. Under certain conditions, the LP relaxation is tight, meaning its solution coincides with that of the original ILP and hence produces the ML codeword in polynomial time. When these conditions fail, however, the LP solution is typically fractional and does not correspond to any valid codeword. In short, this method does not guarantee recovery of the ML codeword in general.
Practical coding standards use particular decoders co-designed with code-specific decoding methods to achieve an ML decoding performance in polynomial time for a restricted class of codes or channels. Prominent examples include Bose–Chaudhuri–Hocquenghem (BCH) codes decoded by the Berlekamp–Massey algorithm [35], decoding of expander codes over the BSC [51], and decoding of first-order Reed-Muller codes [44]. Recently, code-agnostic approaches to ML decoding have attracted significant attention, as they enable universal applicability across arbitrary codes and support practical implementations in reliable communication settings, where the SNR is typically high and moderate-to-high-rate codes are employed. This line of research was initiated by the remarkable work of Duffy et al. [14], who introduced the method of ML decoding via noise guessing, termed GRAND. This framework addresses discrete memoryless additive-noise channels, where the received sequence equals the transmitted codeword plus noise (modulo the alphabet). Because the noise uniquely determines the transmitted codeword given the received sequence, decoding reduces to testing (or, guessing) candidate noise vectors ranked by likelihood. Each is subtracted from the received sequence and checked for codebook membership; the first valid residual yields the ML codeword. The computational cost is governed by the number of guesses required before termination. Although an exact complexity is unknown, it was shown that in the asymptotic regime $n\to\,\infty$, the expected number of guesses scales as $2^{\,n\min\{H_{1/2},\,1-\frac{k}{n}\}}$, where $H_{\alpha}$ denotes the Rényi entropy of order $\alpha$ of the noise (with $H_{1}$ being the Shannon entropy). As a Viterbi decoding approach for linear block codes, this method offers substantially lower complexity than ESD in the high-SNR regime (i.e., when noise entropy is small) and is particularly effective for high-rate, short block codes. Its advantage, however, vanishes as the SNR decreases, the code rate is reduced, and especially when the channel is non-additive.
@@
As a related note, extensions of ML decoding to list decoding, soft-decision decoding (or simply soft decoding), and erasure decoding have been widely studied. For example, Elias’s early work on list decoding (also called as list-$\ell$ decoding, where $\ell$ denotes the list size) [16] showed that allowing the decoder to output multiple candidates (i.e., allow some level of ambiguity) can reduce error probability, while erasure decoding provides a simple way to incorporate partial reliability information. Soft-decision decoding, particularly with ML-based algorithms, has since become the standard in practice [19]. However, each of these approaches is tailored to a specific setting, and no existing method provides a unified, code-agnostic framework that applies simultaneously to ML hard-decision, soft-decision, erasure, and list decoding.
@@ I-B The Proposed Algorithm and its Relation to the Prior Work
This paper introduces a code-agnostic framework that, for the first time, breaks the long-standing $\mathcal{O}(2^{k}n)$ complexity barrier of ESD. The key idea is to express the likelihood of each codeword as the inner product of two vectors: one derived from the received sequence, and the other representing the codeword itself. This allows us to represent the entire codebook as a single, large, binary matrix $\mathbf{M}$. Consequently, ML decoding is reduced to a single vector–matrix multiplication, followed by selecting the maximum entry of the resulting vector.
The efficiency of the proposed approach stems from exploiting the discrete structure of $\mathbf{M}$. This performance gain parallels the classical distinction between sorting arbitrary real numbers and sorting integers. In general, sorting $n$ arbitrary items requires $\Omega(n\log n)$ comparisons (i.e., grows on the order of $n\log n$), whereas algorithms such as Counting Sort—particularly efficient when the range of possible input values is small relative to the number of elements—leverage the finite alphabet of integers to achieve linear $\mathcal{O}(n+k)$ complexity [29], where $k$ denotes the range size of the integer-valued inputs. Likewise, rather than treating the entries of $\mathbf{M}$ as arbitrary reals, we exploit its binary structure via the Mailman algorithm [32], which computes the product of an $m\times n$ binary matrix $\mathbf{A}$ with a real vector $\boldsymbol{x}\in\mathbb{R}^{n}$ in $\mathcal{O}(mn/\log\max\{m,n\})$ amortized time complexity—improving upon the straightforward $\mathcal{O}(mn)$ complexity for general matrices. This contrasts with the classical result of Winograd [49], which establishes that general vector-matrix multiplication requires $\Omega(mn)$ operations. Employing this algorithm reduces the total number of operations to $\mathcal{O}(2^{k})$—an $n$-fold improvement over ESD. For the special case of linear codes over the BSC, the framework further adapts to syndrome decoding with a complexity of $\mathcal{O}(2^{\min\{k,\,n-k\}})$, again improving upon the standard $\mathcal{O}(n\cdot 2^{\min\{k,\,n-k\}})$ complexity of trellis-based ML decoders. Since the time complexity is inherently proportional to the number of codewords (or syndromes, whichever is smaller), no further substantial reduction should be expected—identifying the most likely codeword necessarily entails examining each candidate at least once. The main drawback lies in memory usage: the method stores a matrix representing all codewords in the codebook, resulting in a space complexity of $\mathcal{O}(2^{k})$. Nonetheless, this is still smaller by a factor of $n$ compared to trellis-based (Viterbi) decoding. Moreover, the framework is universal: it applies to both linear and nonlinear block codes, supports hard- and soft-decision decoding, and extends naturally to general memoryless as well as intersymbol-interference (ISI) channels—the most widely studied class of channels with memory [40].
@@
Unlike all previously known ML decoding methods, the proposed decoder universally achieves a complexity of $\mathcal{O}(2^{k})$, which is independent of the block length $n$ when the code rate $k/n$ is assumed to be fixed. This complexity is also independent of the SNR level, distinguishing it from approaches like GRAND [14]. Moreover, the algorithm does not depend on the specific received sequence, unlike the algebraic ML decoder of [27]. The complexity of our proposed method is smaller by a logarithmic factor compared to the best known trellis-based ML decoding via the Viterbi algorithm. This makes the method particularly attractive for low-SNR channels—a relatively unexplored regime where reliable transmission requires low-rate codes with large minimum distance. For example, Reed–Muller (RM) codes $\text{RM}(r,m)$ have long been employed in deep-space and satellite communications due to their high minimum distance and simple majority-logic decoding [36, 38]. For fixed order $r$, the dimension of RM codes scales as $k=\Omega(\log n)$, i.e., it grows logarithmically with the block length $n$. In this regime of very low SNRs and code rates, approximate ML decoding algorithms fail to provide satisfactory performance, and exact ML decoders cannot achieve lower complexity than exhaustive search. Moreover, when extended to list decoding, the proposed algorithm’s complexity increases only logarithmically with the list size, unlike other approaches such as SOGRAND [53] and List Viterbi [45], whose complexity grows linearly with the list size.
@@
The proposed method is especially relevant for emerging applications such as the ultra-reliable low-latency communication (URLLC), a foundational requirement of 5G and future 6G networks. These systems face a fundamental tradeoff: to increase reliability, one often lengthens the block length $n$, which, for conventional universal decoders, directly increases decoding latency, often scaling as $\mathcal{O}\bigl(2^{k}n\bigr)$. In contrast, our method’s $\mathcal{O}(2^{k})$ complexity decouples decoding delay from the block length. There are two key advantages: 1) for a given message/information size $k$, designers can increase $n$ to boost reliability without increasing decoding delay; and 2) the $n$-fold reduction in complexity becomes increasingly significant for large $n$. These advantages make the proposed method a promising enabler for applications that demand both high reliability and ultra-low latency, including Remote surgery & Tactile Internet, Augmented/Virtual/Mixed Reality (AR/VR/XR), Vehicle-to-everything (V2X) communication, and Industrial automation, among many others [31, 34, 54].
@@
Same time complexity as ESD. High space complexity.
@@
Our method. In general $\mathcal{O}(2^{k})$ $\mathcal{O}(2^{k})$ Code-agnostic; Works for all codes for both hard- and
        For linear codes $\mathcal{O}(2^{\min\{k,\,n-k\}})$ $\mathcal{O}(2^{\min\{k,\,n-k\}})$ soft-decision decoding. High space complexity.
@@ I-C Organization
The core contributions of this paper are presented in Section II. Part II-A introduces the universal decoding method for discrete memoryless channels and analyzes its complexity. Part II-B extends the framework to soft-decision decoding and list-$\ell$ ML decoding, and derives the associated complexities. Part II-C applies the method to erasure decoding. Part II-D focuses on linear codes over symmetric channels, showing how syndrome decoding can be used to further reduce complexity. Finally, Part II-E treats the generalization to $L$-tap ISI channels. We conclude the paper in Section IV, where we also discuss the inherent difficulty of the underlying vector-matrix multiplication routine and, drawing on prior work—though without formal proof—argue that further substantial improvements in complexity are unlikely. Finally, a brief overview of the Mailman algorithm is given in Appendix The Mailman Algorithm for Fast vector-matrix Multiplication.
@@ II-A Decoding Algorithm Formulation
which can be accomplished using a trivial linear-time scan: by iterating through the vector once and updating a register that stores the current maximum value. Since the length of the resulting vector $\mathbf{m}$ is $S$, this step has a computational complexity of $\mathcal{O}(S)$. We summarize the procedure in Box II-A, then analyze the overall decoding complexity.
@@
Space complexity. Step 2 of our algorithm employs the Mailman algorithm [32] for fast vector–matrix multiplication by factoring the codebook matrix as $\mathbf{M}=UP$ (see Appendix The Mailman Algorithm for Fast vector-matrix Multiplication). As shown therein, only the permutation matrix $P$, which is a binary matrix of size $S\times S$, needs to be stored at the receiver, and its particular structure allows efficient representation with space complexity $\mathcal{O}(S)=\mathcal{O}(q^{k})$.
@@
The matrix $\mathbf{M}$ and its factorization $\mathbf{M}=UP$ are constructed once at the receiver as an offline pre-processing step. Hence, their computational costs are amortized and excluded from the overall complexity analysis. This treatment follows the convention in prior works, e.g., [55], and parallels the one-time trellis construction required by the Viterbi algorithm [50].